How do u solve the following equation? 3n=-2-4
Found 2 solutions by checkley75, bucky:
Answer by checkley75(3666)

3N=-2-4 COMBINE THE NUMBERS ON THE RIGHT HALF OF THE EQUATION THUS:
3N=-6 NOW DIVIDE BOTH SIDES BY 3
3N/3=-6/3
N=-2 ANSWER.
PROOF
3*-2=-2-4
-6=-6

Answer by bucky(2189)

solve:
.
3n = -2 - 4
.
combine the two numbers on the right side by addition of negative numbers to get:
.
3n = -6
.
Solve for n by dividing both sides by 3. This division makes the equation become:
.
(3/3)*n = (-6/3)
.
which simplifies to:
.
1*n = -2
.
or simply:
.
n = -2
.
Check by substituting this value for n in the original equation that you were given.
This substitution results in:
.
3n = -2 - 4
.
becoming:
.
3*(-2) = -2 - 4
.
the left side multiplies out to:
.
-6 = -2 - 4
.
and the right side adds to -6. So the equation works because it results in:
.
-6 = -6
.
Hope this helps you to understand some of the basic methods of solving linear equations.
.